OpenStudy (annbanz):

log x = log(x^2 - 12)

OpenStudy (anonymous):

Let \(f(x) = e^x\), if you do \(f(x)\) to both sides, you get: \[ f(\log x) = f(\log(x^2-12)) \]\[ x = x^2-12 \]

OpenStudy (annbanz):

Forgot I could cancel out the log's, wow. Thanks @wio!